OpenStudy (anonymous):

Algebra Help Please 1. Solve each equation by graphing the related function. If the equation has no real-number soultion, write no solution. x^2 - 9 = 0 A) x = +3 B) x = +9 C) x = 0 D) no solution

OpenStudy (anonymous):

@DanJS @texaschic101

OpenStudy (danjs):

x^2 will always give a positive value

OpenStudy (anonymous):

@DanJS So it could be any of the answers?

OpenStudy (anonymous):

So it wants you to graph y=x^2-9. This is a parabolic function with it's vertex starting at (0,-9) and growing upwards. Then you need to find where it is equal to 0, ie where the function hits the x-axis. We can do this by setting x^2-9 equal to 0 and solving for x (this will give us two roots, one the opposite sign of the other). Do you understand?

OpenStudy (anonymous):

Obviously, the answers only include one of the two solutions, so you just have to calculate which one :)

OpenStudy (anonymous):

How can I calculate it?

OpenStudy (anonymous):

\[x^2-9=0 \rightarrow x^2=9\] Solve for x.

OpenStudy (anonymous):

9x^2

OpenStudy (anonymous):

I'm guessing B

OpenStudy (anonymous):

Don't guess. Solve it. It's easy.

OpenStudy (anonymous):

I dont understand. Can you talk me threw it?

OpenStudy (anonymous):

\[x^2=9\]We need to solve for x, but we have x^2. We need to fix that. So how do we reduce the power so that we only have x? \[\sqrt{x^2}=\sqrt{9} \rightarrow (x^2)^{1/2}=\sqrt{9}\rightarrow x=??\] What's the square root of 9?

OpenStudy (anonymous):

The square root of 9 = 3

OpenStudy (anonymous):

@CShrix Is the correct answer A ?

OpenStudy (anonymous):

Yes and -3 as well. There are two solutions for x. Now, the question asks us to solve graphically, but we can use the above to check our work :) Okay, so how do we solve this graphically? We can set up a table x | y -3 0 -2 -5 -1 -8 0 -9 1 -8 2 -5 3 0 The roots are anywhere where y=0. We see from our table that this is at y=3 and -3. Checking it again we see that \[(-3)^2-9=0\]and\[(3^2)-9=0\]

OpenStudy (anonymous):

So the answer is really C ?

OpenStudy (anonymous):

` We see from our table that this is at y=3 and -3.` Sorry, I meant where x=3, -3

OpenStudy (anonymous):

Oh okay so the answer is A

OpenStudy (anonymous):

`The roots are anywhere where y=0. We see from our table that this is at x=3 and -3.` :)

OpenStudy (anonymous):

It's D